the complement of an angle is one-third the size of the supplement of the same angle. Find the measure of the angle?
Answer by HyperBrain(694)

Let x=angle
90-x=complement of the angle
180-x=supplement of the angle
---
So, we have the equation:
90-x=1/3(180-x)
multiply both sides by 3, to remove the fraction
3(90-x)=180-x
270-3x=180-x
2x=90
x=45 degrees